A sum of Rs. 500 is in the form of denominations of Rs. 5 and 10. If total no. of notes is 90, find no. of each type of note.

Answer by gokhalesujata(4)

Let x be the number of notes of denomination Rs5
Let y be the number of notes of denomination Rs10
we know that the total amount is Rs500 which means 5 times the number of notes of denom Rs5 + 10 times the number of notes of denomination Rs 10 equals 500
ie. 5x+10y=500
Also total number of notes is 90 ie x+y=90
hence x=90-y
5*(90-y)+10y=500
450-5y+10y=500
5y=500-450
5y=50
y=50/5
y=10
x=90-y=90-10=80
hence 10 notes of Rs10 and 80 notes of Rs5